# Cobordism of Domes Over Curves

### v1

@@ Abstract.
An integral curve is a closed piecewise linear curve comprised of unit intervals. A dome is a polyhedral surface whose faces are equilateral triangles and whose boundary is an integral curve. Glazyrin and Pak showed that not every integral curve can be domed by analyzing the case of unit rhombi, and conjectured that every integral curve is cobordant to a unit rhombus. We show that this is false for oriented domes, but that every integral curve is cobordant to the union of finitely many unit rhombi.
@@ 1. Introduction
Let $\gamma$ be a closed piecewise linear curve in three dimensional Euclidean space $\mathbb{E}$. We say that $\gamma$ is integral if all intervals have integer length. Note that we allow $\gamma$ to have multiple connected components. Now let $S$ be a piecewise linear surface in $\mathbb{E}$ with boundary $\gamma$ and whose facets are all unit triangles. Note that $S$ need not be embedded or immersed. We say that $S$ is a dome over $\gamma$, that $\gamma$ is spanned by $S$, and that $\gamma$ can be domed.
As early as 2005, Kenyon asked if every integral curve can be domed, see  [5]. In 2021, this was shown to be false by Glazyrin and Pak, who proved a necessary condition for a unit rhombus to be domed in  [4]. Moreover, Glazyrin and Pak conjectured that this is in some sense the only restriction that prevents a general integral curve from being domed.
Conjecture 1.1 ([4, Conj.  5.14]).
@@
There exist unit rhombi $\rho_{1},\rho_{2}$ such that $\rho_{1}\cup\rho_{2}$ is not orientably cobordant to any unit rhombus $\rho_{3}$.
Our proof of this negative result is non-constructive, but shows that the statement is true for almost all pairs of unit rhombi $\rho_{1},\rho_{2}$. Our work builds upon techniques introduced by Anan’in and Korshunov, who gave a second proof that Kenyon’s question is false in [1]. Their proof is also non-constructive, and shows that almost all integral curves cannot be domed. They consider a boundary map from the moduli space of domes to the moduli space of integral curves, and prove that its image has measure zero in the orientable case. Our result extends their work by allowing integral curves with multiple components.
We also prove a weaker version of Conjecture 1.1, posed by Glazyrin and Pak in  [4, Conj.  5.15], which allows for finitely many rhombi.
@@
For every integral curve $\gamma$, there is a finite set of unit rhombi $\rho_{1},\dotsc,\rho_{k}$ such that $\gamma$ is cobordant to $\rho_{1}\cup\dotsc\cup\rho_{k}$. Moreover, for $|\gamma|=n$, it suffices to take $k=n^{2}+2n-12$ rhombi.
Our proof is constructive, and uses rhombus pivots to reduce a generic integral curve to a planar integral curve. That is, for consecutive vertices $u,v,w$ in $\gamma$, we can replace $v$ by some point $v^{\prime}\in B_{1}(u)\cap~{}B_{1}(w)$ by attaching the rhombus $[uvwv^{\prime}]$. Then we prove the theorem directly for planar integral curves. Our approach is similar to ideas introduced in [4, §2].
@@
We prove Theorem 1.3 in Section 2 because it is mostly self-contained. Then in Section 3 we introduce necessary notions to describe the moduli space of domes and curves. This is a generalization of many of the definitions and results given in [1] to allow for domes to bound integral curves with multiple connected components. Then we prove Theorem 1.2 in Section 4 with these techniques. Final remarks are given in Section 5.
@@ 2. Every integral curve is cobordant to finitely many unit rhombi
In this section we prove Theorem 1.3, that every integral curve $\gamma\in\mathcal{M}_{n}$ is cobordant to a finite union of rhombi $\rho_{1}\cup\dotsc\cup\rho_{k}$.
We say that two integral curves $\gamma$ and $\eta$ are rhombus equivalent if there exist finitely many unit rhombi $\rho_{1},\dotsc,\rho_{k}$ and a dome over $\gamma\cup\eta\cup\rho_{1}\cup\dotsc\cup\rho_{k}$. Here we say $k$ is the number of rhombi used in the rhombus equivalence. This is similar to the definition of flip equivalence. (See e.g. [4, § 2.4].) In our definition, we do not need to assume that the added rhombi can be domed, which is necessary for flip equivalence. Clearly if an integral curve $\gamma$ is rhombus equivalent to an integral curve which can be domed, then $\gamma$ satisfies the existence condition of Theorem  1.3.
First, we show that every integral curve is rhombus equivalent to a planar integral curve. An integral curve is planar if it lies in a plane $H\subset\mathbb{E}$.
@@
Every integral curve $\gamma$ is rhombus equivalent to a planar integral curve. Moreover, for $|\gamma|=n$, it suffices to use $k={n\choose 2}$ rhombi.
@@
Choose two vertices $v$ and $w$ of $\gamma$ that are a maximum distance apart, and take any plane $H$ containing the line $(v,w)$. This gives a decomposition of $\gamma$ as two integral paths containing $v$ and $w$, and we show that both can be made to lie entirely in $H$ via rhombus equivalence.
Suppose $\eta$ is an integral path with vertices $v=v_{1},\dotsc,v_{m},v_{m+1}=w$. Let $h_{i}$ denote the distance from $v_{i}$ to the plane $H$. (See Figure 1.) Choose $v_{i}$ maximizing $h_{i}$. If the same height is achieved by multiple  $v_{i}$, choose the one with the smallest index $i$. Note that this means $h_{i-1}<h_{i}\geq h_{i+1}$.
@@
Consider the circle $B_{1}(v_{i-1})\cap B_{1}(v_{i+1})$ which contains $v_{i}$. For each point $v_{i}^{\prime}\in B_{1}(v_{i-1})\cap B_{1}(v_{i+1})$, we can replace $v_{i}$ by $v_{i}^{\prime}$ in $\eta$ by rhombus equivalence via $[v_{i-1}v_{i}v_{i+1}v_{i}^{\prime}]$. (See Figure 2.) Choose $v_{i}^{\prime}$ to minimize $h_{i}^{\prime}$, the new distance from $v_{i}^{\prime}$ to $H$. If $h_{i}^{\prime}=0$, then $v_{i}^{\prime}$ is in $H$. Otherwise if $h_{i}^{\prime}>0$, minimizing $h_{i}^{\prime}$ corresponds to choosing $v_{i}^{\prime}$ so that the slope of the line $(v_{i-1},v_{i}^{\prime})$ is parallel to the slope of the line $(v_{i},v_{i+1})$. Because $h_{i+1}\leq h_{i}$, we see that $h_{i}^{\prime}\leq h_{i-1}<h_{i}$, so we can always decrease the maximum height $h_{i}$ by a rhombus equivalence.
@@
Each added rhombus will remove at least one inversion from the sequence $h_{1},\dotsc,h_{m}$. I.e., $h_{i-1}<h_{i}$ is replaced with $h_{i-1}\geq h_{i}^{\prime}$. Since there can be at most (${m\choose 2}$ inversions, after at most (${m\choose 2}$ flips we guarantee the integral path $\eta$ lies in $H$. Therefore, for the whole integral curve $\gamma$, we will need at most (${m\choose 2}+{n-m\choose 2}\leq{n\choose 2}$ flips. ∎
Now we prove the result of Theorem 1.3 for planar integral curves. Our proof is by induction on $n=|\gamma|$, and divides the integral curves into smaller pieces. We say that an integral curve $\gamma=[v_{1}\dotsc v_{n}]$ is $\epsilon$-packing around $v_{i}$ if for all $j$, the length $|v_{i},v_{j}|<\epsilon$. In this case, we say that $v_{i}$ is the packing center of $\gamma$, and when the packing center is understood we just say that $\gamma$ is $\epsilon$-packing.
@@
Every planar integral curve $\gamma$ is rhombus equivalent to a planar integral curve which is $2$-packing. Moreover, if $|\gamma|=n$, then it suffices to use $k={n\choose 2}$ rhombi.
@@
The existence of a rhombus equivalent $2$-packing curve follows from the Steinitz Lemma, which can be stated as follows. (See e.g.  [2, Theorem  2.1].) For each dimension $d>0$, there is a constant $B_{d}$ such that for unit vectors $u_{1},\dotsc,u_{n}\in\mathbb{R}^{d}$ satisfying $u_{1}+\dotsc+u_{n}=0$, there is a permutation $\sigma\in S_{n}$ such that for each $1\leq i\leq n$,
$|u_{\sigma(1)}+\dotsc+u_{\sigma(i)}|\leq B_{d}$
For $d=2$, Bergström showed the optimal value is $B_{2}=\sqrt{5}/2$ in [3]. However, we use the constant $2$ because it is sufficient for the proof of Theorem 1.3 and gives clearer exposition.
Viewing the integral curve $\gamma=[v_{1}\dotsc v_{n}]$ as a set of vectors $u_{1},\dotsc,u_{n}$, with $u_{i}$ pointing from $v_{i}$ to $v_{i+1}$, we can swap a pair of consecutive vectors $u_{i}$ and $u_{i+1}$. By adding a rhombus containing the endpoints $v_{i-1},v_{i},v_{i+1}$, and the point $v_{i}$ reflected across the line $(v_{i-1},v_{i+1})$. This corresponds to the simple transposition $(i,i+1)\in S_{n}$. And because we can achieve any permutation $\sigma\in S_{n}$ by the product of simple transpositions, there exists a rhombus equivalent $2$-packing planar integral curve around $v_{1}$. (See e.g. [6].) Moreover, for every permutation $\sigma\in S_{n}$, the maximal length of $\sigma$ in terms of simple transpositions is (${n\choose 2}$, so this number of rhombi is sufficient. ∎
Now we are ready to prove that every integral curve is cobordant to finitely many unit rhombi.
@@
Every integral curve $\gamma$ is cobordant to $\rho_{1}\cup\dotsc\cup\rho_{k}$ for a finite set of unit rhombi $\rho_{1},\dotsc,\rho_{k}\in\mathcal{M}_{4}$. Moreover, for $|\gamma|=n$, it suffices to take $k=n^{2}+2n-12$ rhombi.
@@
From Lemma 2.1 and Lemma 2.2, we may assume without loss of generality that $\gamma$ is planar and $2$-packing. This uses $n(n-1)$ rhombi. Our proof is by induction. The base case is when $|\gamma|=5$.
Suppose $\gamma=[v_{1}\dotsc v_{5}]$. If the circumradius of the triangle $[v_{1}v_{3}v_{4}]$ is less than $1$, then there is a point $z\in B_{1}(v_{1})\cap B_{1}(v_{3})\cap B_{1}(v_{4})$. (See Figure 3, left.) Adding the edges $[z,v_{1}],[z,v_{3}]$ and $[z,v_{4}]$ gives a cobodism between $\gamma$ and two unit rhombi $[v_{1}v_{2}v_{3}z]$ and $[v_{1}zv_{4}v_{5}]$. If the circumradius of the triangle $[v_{1}v_{3}v_{4}]$ is greater than $1$, then we can perform a rhombus equivalence to construct a new integral curve $\gamma^{\prime}$ with the desired circumradius. Note that at most $3$ rhombi are used.
Figure 3. The base case and inductive step.
In the inductive step, for a planar, $2$-packing integral curve $\gamma=[v_{1}\dotsc v_{n}]$ with $n>5$, there is some point $z\in B_{1}(v_{1})\cap B_{1}(v_{4})\cap H$, where $H$ is the plane containing $\gamma$. (See Figure 3, right.) Now $[v_{1}v_{2}v_{3}v_{4}z]$ is a planar pentagon, so the base case applies, and $[v_{1}zv_{4}\dotsc v_{n}]$ is a planar, $2$-packing integral curve of length $n-1$, so the inductive hypothesis applies. This divides $\gamma$ into $n-4$ pentagons, each of which use at most $3$ rhombi. So in total at most $n(n-1)+3(n-4)=n^{2}+2n-12$ rhombi are used. ∎

### v2

@@ Abstract.
An integral curve is a closed piecewise linear curve comprised of unit intervals. A dome is a polyhedral surface whose faces are equilateral triangles and whose boundary is an integral curve. Glazyrin and Pak showed that not every integral curve can be domed by analyzing the case of unit rhombi, and conjectured that every integral curve is cobordant to a unit rhombus. We show that this is false for oriented domes, but that every integral curve is orientably cobordant to the union of finitely many unit rhombi.
@@ 1. Introduction
Let $\gamma$ be a closed piecewise linear curve in three dimensional Euclidean space $\mathbb{E}$. We say that $\gamma$ is integral if all intervals have unit length. We also consider the case where $\gamma$ has multiple connected components, in which case we say that $\gamma$ is a union of integral curves. Now let $S$ be a piecewise linear complex in $\mathbb{E}$ with boundary $\gamma$ and whose facets are all unit triangles. Note that $S$ need not be embedded or immersed. We say that $S$ is a dome over $\gamma$, that $\gamma$ is spanned by $S$, and that $\gamma$ can be domed.
As early as 2005, Kenyon asked if every integral curve can be domed, see  [Ken05]. In 2022, this was shown to be false by Glazyrin and Pak, who proved a necessary condition for a unit rhombus to be domed in  [GP22]. Moreover, Glazyrin and Pak conjectured that this is in some sense the only restriction that prevents a general integral curve from being domed.
Conjecture 1.1 ([GP22, Conj.  5.14]).
@@
There exists an integral curve $\gamma$ of length $5$ which is not orientably cobordant to any unit rhombus $\rho$.
Our proof of this negative result is non-constructive. We show that for almost all pairs of unit rhombi $\rho_{1},\rho_{2}$, the union $\rho_{1}\cup\rho_{2}$ is not orientably cobordant to any third unit rhombus $\rho_{3}$. Then we give a cobordism between $\rho_{1}\cup\rho_{2}$ and $\gamma$. Our work builds upon techniques introduced by Anan’in and Korshunov, who gave a second proof that Kenyon’s question is false in [AK24]. Their proof is also non-constructive, and shows that almost all integral curves cannot be domed. They consider a boundary map from the moduli space of domes to the moduli space of integral curves, and prove that its image has measure zero in the orientable case. Our result extends their work by allowing unions of integral curves.
We also prove a weaker version of Conjecture 1.1, posed by Glazyrin and Pak in  [GP22, Conj.  5.15], which allows for finitely many rhombi.
@@
For every integral curve $\gamma$, there is a finite set of unit rhombi $\rho_{1},\dotsc,\rho_{k}$ such that $\gamma$ is orientably cobordant to $\rho_{1}\cup\dotsc\cup\rho_{k}$. Moreover, for $|\gamma|=n$, it suffices to take $k=n^{2}+2n-12$ rhombi.
Our proof is constructive, and uses rhombus equivalence to reduce a generic integral curve to a planar integral curve. We use $B_{r}(v)$ to mean the sphere of radius $r>0$ around a point $v\in\mathbb{E}$, and for a set of points $v_{1},\dotsc,v_{n}\in E$, we use $[v_{1}\dotsc v_{n}]$ to mean the integral curve through $v_{1},\dotsc,v_{n},v_{n+1}=v_{1}$. For consecutive vertices $u,v,w$ in $\gamma$, we can replace $v$ by some point $v^{\prime}\in B_{1}(u)\cap~{}B_{1}(w)$ by attaching the rhombus $[uvwv^{\prime}]$. We call this a rhombus equivalence. Then we prove the theorem directly for planar integral curves. Our approach is similar to ideas introduced in [GP22, §2].
@@
We prove Theorem 1.3 in Section 2 because it is mostly self-contained. Then in Section 3 we introduce necessary notions to describe the moduli space of domes and curves. This is a generalization of many of the definitions and results given in [AK24] to allow for domes to bound unions of integral curves. Then we prove Theorem 4.3 in Section 4 with these techniques. Final remarks are given in Section 5.
@@ 2. Every integral curve is cobordant to finitely many unit rhombi
In this section we prove Theorem 1.3, that every integral curve $\gamma\in\mathcal{M}_{n}$ is orientably cobordant to a finite union of rhombi $\rho_{1}\cup\dotsc\cup\rho_{k}$.
We say that two integral curves $\gamma$ and $\eta$ are rhombus equivalent if there exist finitely many unit rhombi $\rho_{1},\dotsc,\rho_{k}$ and a dome over $\gamma\cup\eta\cup\rho_{1}\cup\dotsc\cup\rho_{k}$. Here we say $k$ is the number of rhombi used in the rhombus equivalence. This is similar to the definition of flip equivalence. (See e.g. [GP22, § 2.4].) We also want to distinguish when a rhombus equivalence is orientable. However, when some of the integral curves share common edges, the resulting dome will not be purely $2$-dimensional, so we define orientability carefully. For an integral curve $\gamma$, we call the surface formed by gluing the boundary of a disk to $\gamma$ the closure of $\gamma$. We say a dome over $\gamma,\eta,\rho_{1},\dotsc,\rho_{k}$ is orientable if the dome, along with the closure of $\gamma,\eta,\rho_{1},\dotsc,\rho_{k}$ forms a closed, orientable surface. Clearly, if an integral curve $\gamma$ is orientably rhombus equivalent to an integral curve which can be domed, then $\gamma$ satisfies the existence condition of Theorem  1.3, and we can prove the theorem by keeping track of the number of rhombi used.
First, we show that every integral curve is orientably rhombus equivalent to a planar integral curve. An integral curve is planar if it lies in a plane $H\subset\mathbb{E}$.
@@
Every integral curve $\gamma$ is orientably rhombus equivalent to a planar integral curve. Moreover, for $|\gamma|=n$, it suffices to use $k={n\choose 2}$ rhombi.
@@
Choose two vertices $v$ and $w$ of $\gamma$, and take any plane $H$ containing the line $(v,w)$. This gives a decomposition of $\gamma$ as two integral paths containing $v$ and $w$, and we show that both can be made to lie entirely in $H$ via rhombus equivalence.
Suppose $\eta$ is an integral path with vertices $v=v_{1},\dotsc,v_{m},v_{m+1}=w$. Let $h_{i}$ denote the distance from $v_{i}$ to the plane $H$, hence $h_{1}=h_{m+1}=0$. (See Figure 1.) Note that the heights $h_{i}$ are all nonnegative, but the $v_{i}$ may lie on different sides of $H$; this will not affect the proof. Choose $v_{i}$ maximizing $h_{i}$. If the same height is achieved by multiple  $v_{i}$, choose the one with the smallest index $i$. Note that this means $h_{i-1}<h_{i}\geq h_{i+1}$.
@@
Consider the circle $B_{1}(v_{i-1})\cap B_{1}(v_{i+1})$ which contains $v_{i}$. For each point $v_{i}^{\prime}\in B_{1}(v_{i-1})\cap B_{1}(v_{i+1})$, we can replace $v_{i}$ by $v_{i}^{\prime}$ in $\eta$ by rhombus equivalence via $[v_{i-1}v_{i}v_{i+1}v_{i}^{\prime}]$. (See Figure 2.) Choose $v_{i}^{\prime}$ to minimize $h_{i}^{\prime}$, the new distance from $v_{i}^{\prime}$ to $H$. If $h_{i}^{\prime}=0$, then $v_{i}^{\prime}$ is in $H$ and we are done with this vertex. If not, then we can at least guarantee that $h_{i}^{\prime}<h_{i}$. Choosing $v_{i}^{\prime}$ so that the line $(v_{i-1},v_{i}^{\prime})$ is parallel to the line $(v_{i},v_{i+1})$, (i.e., choosing $v_{i}^{\prime}$ so that the rhombus $[v_{i-1}v_{i}v_{i+1}v_{i}^{\prime}]$ is planar) we see that $h_{i}^{\prime}\leq h_{i-1}<h_{i}$ because $h_{i+1}\leq h_{i}$. Thus we can always decrease the maximum height $h_{i}$ by a rhombus equivalence.
Note also that either $v_{i-1}$ or $v_{i+1}$ may be on the other side of $H$ relative to $v_{i}$, in which case the circle $B_{1}(v_{i-1})\cap B_{1}(v_{i+1})$ will always intersect $H$. (If $B_{1}(v_{i-1})\cap B_{1}(v_{i+1})$ did not intersect $H$, then $h_{i}$ would not be a maximum height.) So in this case, we can always choose $v_{i}^{\prime}\in H$.
@@
To prove our result, we view rhombus equivalence as an operation on the sequence $00,h_{2},\dotsc,h_{m},$ 0 which, in the worst case, takes the first maximum element $h_{i}$ and replaces it with the smaller value of $h_{i-1}$. Because $h_{1}=0$, after repeated application, rhombus equivalence will eventually reduce the height sequence to all 00s. If the maximum height is achieved $k$ separate times, then it will take $k$ rhombus equivalences to reduce the maximum height of the sequence to a smaller value. Thus the worst case is when $h_{1},\dotsc,h_{m}$ are all distinct, when (${m\choose 2}$ flips are needed. So for the whole integral curve $\gamma$, we will need at most (${m\choose 2}+{n-m\choose 2}\leq{n\choose 2}$ rhombi to find an integral curve which lies entirely in $H$.
To see that the rhombus equivalence is orientable, note that the closure of all the rhombi used in the rhombus equivalence is homeomorphic to an annulus. Thus the cobordism over $\gamma$, all of the rhombi used, and the resulting planar integral curve is orientable. ∎
We say that an integral curve $\gamma=[v_{1}\dotsc v_{n}]$ is $\epsilon$-packing around $v_{i}$ if for all $j$, the length $|v_{i},v_{j}|<\epsilon$. In this case, we say that $v_{i}$ is the packing center of $\gamma$, and when the packing center is understood we just say that $\gamma$ is $\epsilon$-packing. Next we prove that every planar integral curve is orientably rhombus equivalent to a $2$-packing planar integral curve.
@@
Every planar integral curve $\gamma$ is orientably rhombus equivalent to a planar integral curve which is $2$-packing around one of its vertices. Moreover, if $|\gamma|=n$, then it suffices to use $k={n\choose 2}$ rhombi.
@@
The existence of a rhombus equivalent $2$-packing curve follows from the Steinitz Lemma, which can be stated as follows. (See e.g.  [Bar08, Theorem  2.1].) For each dimension $d>0$, there is a constant $\beta_{d}$ such that for any unit vectors $u_{1},\dotsc,u_{n}\in\mathbb{R}^{d}$ satisfying $u_{1}+\dotsc+u_{n}=0$, there is a permutation $\sigma\in S_{n}$ such that for each $1\leq i\leq n$,
$|u_{\sigma(1)}+\dotsc+u_{\sigma(i)}|\leq\beta_{d}$
For $d=2$, Bergström showed the optimal value is $\beta_{2}=\sqrt{5}/2$ in [Ber31]. However, we use the constant $2$ because it is sufficient for the proof of Theorem 1.3 and gives clearer exposition.
Viewing the integral curve $\gamma=[v_{1}\dotsc v_{n}]$ as a set of vectors $u_{1},\dotsc,u_{n}$, with $u_{i}$ pointing from $v_{i}$ to $v_{i+1}$, we can swap a pair of consecutive vectors $u_{i}$ and $u_{i+1}$. We add a rhombus containing the endpoints $v_{i-1},v_{i},v_{i+1}$, and the point $v_{i}$ reflected across the line $(v_{i-1},v_{i+1})$. This corresponds to the simple transposition $(i,i+1)\in S_{n}$. And because we can achieve any permutation $\sigma\in S_{n}$ by the product of simple transpositions, there exists a rhombus equivalent $2$-packing planar integral curve around $v_{1}$. (See e.g. [Sta12].) Moreover, for every permutation $\sigma\in S_{n}$, the maximal length of $\sigma$ in terms of simple transpositions is (${n\choose 2}$, so this number of rhombi is sufficient.
Again, we see that this rhombus equivalence is orientable because the closure of all of the rhombi used in the rhombus equivalence is homeomorphic to an annulus. ∎
Now we prove the result of Theorem 1.3 for planar integral curves. Our proof is direct when $|\gamma|=5$, and when $|\gamma|>5$ we divide $\gamma$ into pentagons.
@@
Every integral curve $\gamma$ is orientably cobordant to $\rho_{1}\cup\dotsc\cup\rho_{k}$ for a finite set of unit rhombi $\rho_{1},\dotsc,\rho_{k}\in\mathcal{M}_{4}$. Moreover, for $|\gamma|=n$, it suffices to take $k=n^{2}+2n-12$ rhombi.
@@
From Lemma 2.1 and Lemma 2.2, we may assume without loss of generality that $\gamma$ is planar and $2$-packing around one of its vertices $v_{1}$. This uses $n(n-1)$ rhombi. Now we prove the statement for planar, $2$-packing integral curves by dividing them into pentagons.
First, suppose $\gamma$ is a pentagon $[v_{1}\dotsc v_{5}]$. If the circumradius of the triangle $[v_{1}v_{3}v_{4}]$ is less than $1$, then there is a point $z\in B_{1}(v_{1})\cap B_{1}(v_{3})\cap B_{1}(v_{4})$. (See Figure 3, left.) Adding the edges $[z,v_{1}],[z,v_{3}]$ and $[z,v_{4}]$ gives a cobordism between $\gamma$ and two unit rhombi $[v_{1}v_{2}v_{3}z]$ and $[v_{1}zv_{4}v_{5}]$. If the circumradius of the triangle $[v_{1}v_{3}v_{4}]$ is greater than $1$, then we can perform a rhombus equivalence to construct a new integral curve $\gamma^{\prime}$ with the desired circumradius. Note that at most $3$ rhombi are used.
Figure 3. The $n=5$ case and the general case.
In the general case, for a planar, $2$-packing integral curve $\gamma=[v_{1}\dotsc v_{n}]$ with $n>5$, there is some point $z\in B_{1}(v_{1})\cap B_{1}(v_{4})\cap H$, where $H$ is the plane containing $\gamma$. (See Figure 3, right.) We can divide $\gamma$ into the planar pentagon $[v_{1}v_{2}v_{3}v_{4}z]$ and $\gamma^{\prime}:=[v_{1}zv_{4}\dotsc v_{n}]$, which is a planar, $2$-packing integral curve of length $n-1$. Continuing this procedure divides $\gamma$ into $n-4$ pentagons, each of which use at most $3$ rhombi. So in total at most $n(n-1)+3(n-4)=n^{2}+2n-12$ rhombi are used.
To see that this cobordism is orientable, note that when $n=5$ the surface formed from the triangle and the closure of the rhombi used is homeomorphic to a disk. And in the general case, we just divide $\gamma$ into the union of pentagons. ∎